OpenStudy (anonymous):

graph 4x^2+y^2=9 what are its lines of symmetry?

OpenStudy (anonymous):

In a circle, you must have x^2 and y^2. The signs must be positive and the coefficients must be the same.

OpenStudy (anonymous):

In a parabola, you must have x and y^2 or y and x^2. Then you know the graph is a parabola.

OpenStudy (anonymous):

n a hyperbola, you must have x^2 and y^2 but one of them is negative.

OpenStudy (anonymous):

these are the answer choices. im still a little confused.

OpenStudy (anonymous):

sorry power went out

OpenStudy (anonymous):

So now we know that 4x^2 + y^2 = 9 is an ellipse because we have x^2 and y^2. Both are positive and the coefficients are different.

OpenStudy (anonymous):

So now your job is to rearrange it so it is in this form: 1 sec

OpenStudy (anonymous):

\[\frac{ x^2 }{ a^2 }+\frac{ y^2 }{ b^2 }=1\]

OpenStudy (anonymous):

how do I rearrange them?

OpenStudy (anonymous):

Do you notice that the right side has to be 1?

OpenStudy (anonymous):

you're saying the same things that the one guy was saying for the other question -,- its almost as if you're copying & pasting. I really need help asap.

OpenStudy (anonymous):

exaclty he ex[plains it well

OpenStudy (anonymous):

so what is the answer?

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

Thank you! can you help me with another one?